# Computational lexical analysis of Flamenco genres

### v1

@@ I Introduction
Flamenco, a rich oral musical tradition originating in Andalusia, a region in southern Spain, reflects a diverse tapestry of cultural influences. Throughout the centuries, this region, along with its musical expression, has been significantly shaped by various cultural settlements. Notably, the influences of Jews, Muslims, and folkloric music from different Spanish regions are discernible contributors, yet undeniably, the cultural imprint of Andalusian gipsies is deeply embedded in the essence of Flamenco ([14], p. 469-473). Recognizing its distinct features and profound significance for the cultural identity of Andalusia, Flamenco was honored with inclusion in the UNESCO List of Intangible Cultural Heritage of Humanity in 2010 [15].
@@
Flamenco, as an oral music tradition, entails the transmission of songs and musical nuances across generations. Consequently, written scores are a rarity, given the intricate and complex melodic ornamentation inherent in vocal melodies. Even the records of renowned Flamenco festivals lack information about some features of the performances, including the lyrics [20]. The process of manual annotation, essential for capturing these nuances, is not only exceedingly time-consuming but also inevitably involves a subjective layer of interpretation.
@@
To massively analyze large bunches of text we dispose of Natural Language Processing (NLP) tools. The newborn field of Lyrics Information Processing (LIP) [24] has proven to be successful in retrieving all kinds of information through the use of NLP for a computational analysis of lyrics outside Flamenco. Related works on automatic analysis of lyrics cover a wide range of applications such as the author classification of German songs [25], automatic detection of sexism in English songs [26], the age classification of Indonesian lyrics [27] or the identification of different writing styles in Taylor Swift’s songs [28], among others [29, 30, 31, 32]. However, to our knowledge, the computational analysis of Flamenco singers’ production is limited to studies involving analysis of melody and related features [33, 34] rather than the content of their lyrics. We thus face an absence of LIP applications for Flamenco lyrics analysis.
@@
    Can Flamenco palos be classified solely by means of their lexicon? If so, what is the characteristic lexicon of each palo?
@@ II.1 Dataset
We analyze a corpus of Flamenco lyrics curated by Norman P. Kliman on his website [23]. This corpus agglutinates 3380 lyrics from 78 different styles or palos, featuring 276822 tokens (words) and 12981 types (unique words). The corpus is varied and diverse, exhibiting a structure characteristic of natural language corpora (see Appendix A).
@@
Also, for the sake of clarity while reading this manuscript, we specify in Table 1 the acronyms with which we will refer to the aforementioned styles throughout the text and in the Figures. It is important to note that categorizing Flamenco songs into styles can be subjective, but for validation purposes, we rely on the classification provided by the album in which each song appears.
@@
After preprocessing the corpus, a necessary step previous to their computational analysis which involves deleting punctuation signs and accents and distinguishing between common and proper nouns (see Appendix A), we can extract information from the lexical structure of the lyrics. We now proceed to describe several findings in the process of corpus characterization.
@@
One way to quantify the lexical richness of a text is to measure its vocabulary size $|V|$, i.e., the amount of types. In Fig. 2 we show the vocabulary size and the amount of tokens of the eight different styles or palos, computed as the vocabulary size of an aggregation of all the lyrics of each palo. Bulerías account for the highest number of types used while malagueñas restrict the range of use to a vocabulary almost 6 times lower in size. The overall ranking of token and type frequencies remains consistent across genres, with the exception of seguiriyas.
Figure 2: Amount of tokens $L$ (bars on the right) and types $|V|$ (bars on the left) for each style.
In the context of song-wise analysis, we may also observe in Fig. 3 that bulerías have the widest distribution of types used among songs, while malagueñas, fandangos and seguiriyas are restricted to a similar number of types among songs. Cases like malagueñas or alegrías resemble bimodal distributions, indicating that mainly two different vocabulary sizes characterize their lyrics. With these remarks, we can already infer that bulerías serve as the most polyvalent style, with the broadest vocabulary size distribution and the highest amount of types.
Figure 3: Distribution of the vocabulary size, $|V|$ (number of types) for each style. The continuous distribution is derived from a discrete histogram using the kernel density estimation method [36].
@@
In our case, Eq. 1 is affected by the fact that some genres use more types than others, as we see in Fig. 2, while also having different amount of tokens and lyrics. For that reason, we use the standardized $TTR$, $sTTR$, which is defined as the $TTR$ averaged over several portions of the same document, with all the portions having the same number of tokens. For each palo we create a document compiling all the genre’s lyrics. Malagueñas is the genre with the least amount of tokens. Then, we take 50 portions of each genre with the length of the entire malagueñas subcorpus and compute the mean $sTTR$ and its error except for malagueñas, for which we take the entire genre a single time without error. We have checked that increasing the portion number above 50 does not significantly alter our results. We show the results in Fig. 4, where the solid line accounts for the $sTTR$ of the entire corpus, our null model for comparison.
@@
The different values of $sTTR$ provide a partial answer to RQ1 by revealing that various Flamenco genres are distinguished by their unique vocabulary distributions and levels of richness. However, it is not clear that these measurements would allow for an unambiguous identification of a palo given some lyrics. For example, the vocabulary distribution is heavily influenced by the specific corpus used, and the reliability of $sTTR$ as a measure of lexical diversity independent of text length has been challenged [38]. Moreover, $sTTR$ values for some palos are close enough to be indistinguishable, as in the case of tangos and tientos, or bulerías and alegrías. To provide a comprehensive answer to RQ1, we need a more sophisticated approach. Thus, we will now examine the specific lexicon utilized by each palo using a machine learning algorithm which has proven to be useful in text analysis.
@@ II.2 Palo classification
We employ a Multinomial Naive Bayes (MNB) model [35], a machine learning classification algorithm, to classify lyrics into distinct Flamenco genres or palos solely based on their lexical content. Initially, the corpus is partitioned into a training set comprising 85% of the data and a validation set containing the remaining 15%. The algorithm is then trained using the training set, leveraging the palo classifications provided in the dataset. Subsequently, the model performance is assessed using the validation set, wherein the genres assigned by the model are compared against the true genre labels associated with the lyrics. Further details about how this classification algorithm works are included in Appendix B.1.
@@
Quite generally, different arrangements of lyrics for the training and validation sets yield different results. For this reason, we compute the accuracy of the model (the ratio of correct assignments) for 100 different trainings, obtaining the results shown in Fig. 6.
Figure 6: Violin plot of the distribution of accuracies obtained for each genre across 100 trainings of the Multinomial Naive Bayes model using various corpus splittings for training and validation.
Despite the fact that genres like soleá (85% mean accuracy), alegrías (88%) and seguiriyas (91%) are classified with outstanding accuracies, a few lyrics are incorrectly classified and therefore confused by the model. To examine these discrepancies, in Fig. 7 we plot the average confusion matrix of 100 different trainings and account only for the incorrect assignments.
@@
Figure 7: Average confusion matrix of the MNB for 100 different trainings, showing the percentage of confusions of each Flamenco style with other genres. The normalization is based only on the confusions.
It is evident how the aforementioned confusions affect the accuracy of the model, as in the cases of tangos (mean accuracy of 50%) and tientos (61%). However, Fig. 6 shows that soleá, alegrías and seguiriyas are classified with outstanding accuracies. Thus, we can assert that a classification of Flamenco lyrics based solely on their lexicon is quantitatively possible (RQ1), and that the inaccuracies are due to interesting palo relationships. Now, let us extract the lexical features that allow the model to classify the lyrics, i.e., the characteristic lexicon of each style.
@@ II.3 Characteristic lexicon extraction
The MNB computes a metric called log-probability, $\log P(w|C)$, of each word $w$ for each palo $C$ (see Appendix B.1). The higher $P(w|C)$ is, the more characteristic $w$ is for $C$. We compute $P(w|C)$ for several trainings of the algorithm and obtain results as shown in Fig. 8, where we can see the 15 most characteristic words, i.e., the 15 words with the largest value of $\log P(w|C)$, for two example styles, $C=\mathrm{A}$ (alegrías) [Fig. 8a)] and $C=\mathrm{Se}$ (seguiriyas) [Fig. 8c)].
Figure 8: Summary of the most characteristic words for alegrías and seguiriyas. We plot the log-probability of the 15 most characteristoc words for a) alegrías and c) seguiriyas. Additionally, we show two wordcloud diagrams in which the words’ sizes reflect how characteristic they are for b) alegrías and d) seguiriyas.
While alegrías feature as most characteristic words references to places such as Cadiz, Muralla Real ’Royal Wall’ or mar ’sea’ (all related to Cadiz and its harbors), and other less specific words such as llevar ’to take/bring’, voy ’I go’, conmigo ’with me’, seguiriyas are biased towards abstract and profound concepts such as Dios ’God’, alma ’soul’, libertad ’freedom’ and desgracia ’misfortune’. We will later interpret the characteristic words in Sec. III, and see how these words clearly differentiate between genres.
@@
The number of essential words within each palo serves for distinguishing among various Flamenco genres. Table 2 presents the count of essential words per palo, denoted as $N_{e}$. Notably, while malagueñas have 16 essential words, soleá’s essential lexicon is significantly larger, averaging 186 words, i.e., nearly 12 times more. Despite this, as illustrated in Fig. 2, malagueñas exhibit a far smaller vocabulary compared to soleá. To account for the impact of vocabulary size on the essential lexicon, we normalize $N_{e}$ by the vocabulary size, $|V|$, as depicted in Fig. 10. When adjusted for vocabulary size, seguiriyas boasts the largest essential lexicon, while soleá’s normalized essential lexicon (5.5%) is only four times greater than that of malagueñas (1.5%). The complete lists of these words for all Flamenco genres in descending order of log-probability are included in Appendix C. We emphasize that the relevance of these words is automatically detected and is therefore free of subjective biases. This is a clear advantage of our application of computational classification methods.
@@ II.4 Relationship between palos
The distances between Flamenco styles are therefore helpful in addressing RQ2. To provide a more concrete analysis, we adopt a network representation. In Fig. 12, we present a fully-connected network illustrating the distances between styles, enabling a straightforward visual assessment of their relationships. Each node represents a style, and the edges vary in thickness and color depending on the cosine distance values. The network structure provides insights into the relationships among different palos in several ways. As observed in the dendrogram, alegrías emerges as the most lexically distinct palo. The remaining palos form a more tightly interconnected network, with bulerías, soleá, and tangos prominently contributing to its connectivity. Additionally, the size of the nodes in the network graph corresponds to their closeness centrality, a measure in network analysis that quantifies how close a node is to all other nodes in the network ([40], p. 146), suggesting their significance in facilitating connections between different musical forms. This network representation proves advantageous as it facilitates the extraction of spanning trees ([40], p. 27), which are sub-networks that connect all genres such that each genre is linked directly or indirectly to every other genre, without creating any cycles, thereby maintaining the network connectivity. By selecting the minimum spanning tree (MST), which minimizes the sum of distances between palos, we obtain the representation depicted in Fig. 13, i.e., the MST of Flamenco palos. This approach enables a structured exploration of the relationships between palos, shedding light on their hierarchical organization within the Flamenco genre.
Figure 12: Network relationships between different styles or palos, with each node representing a palo or style. The labels on the nodes follow the acronyms of Table 1. The thickness and color of the edges indicate the distances between the palos, and the size of each node is proportional to its closeness centrality.
Figure 13: Minimum spanning tree of Flamenco palos, showing four branches connected by bulerías. The labels on the nodes follow the acronyms of Table 1.
@@ III Discussion
At first glance, regarding the $sTTR$ analysis, in Fig. 4 we can see that $sTTR$ is genre-distinctive. Indeed, returning to the results of the vocabulary size distribution in Fig. 3, if we focus on the $sTTR$ values, we can see how bulerías make lower use of their types than malagueñas do, even though the vocabulary of the latter is significantly lower than the vocabulary of the former. This indicates that while bulerías are rather varied and move within a wide range of topics, this genre makes relatively less use of their topics than malagueñas do. Values above the null model’s $sTTR$ indicate a more developed use of a specific set of types, while values below indicate either a wider use of types or a less developed use of the available types.
@@
First, we find words related to geographical places and day-to-day life. As previously mentioned, Flamenco lyrics are transmitted orally and encompass a tradition of livings and history. Thus it is common to find references to these topics. Thus, while alegrías feature lexicon related to Cádiz, Navarre and titirimundis ’puppets show’, malagueñas refer to the Corte ’Court, Capital’ (Madrid), tangos talk about wars and bulerías feature references to Seville and proper names like Amparo, Dolores or María.
Secondly, we also find references to gipsies and their families and culture, due to the intense influence of this ethnic group in Flamenco art throughout its history. We obtain specific words like gitano/a ’gipsy’, and common words like madre ’mother’, padre ’father’, or primo/a ’cousin’, usually in a gipsy context. Interestingly, seguiriyas and soleá genres even feature Caló vocabulary [21] as characteristic words. The use of Caló, a dialect of Romani language spoken by the gipsies of Spain ([41], p. 125), further connects these genres to the gipsy people.
We find evidence for the belief that Flamenco accompanied gipsies and Andalusian people in times of sorrow and high suffering [22] in lexicon related to deep pain, such as dolor ’pain’, desgracia ’misfortune’, llorar ’to cry’, pena ’sorrow’, traición ’betrayal’, muerta ’dead (singular feminine)’. This semantic field constitutes a third group, along with lexicon related to joy and love. Flamenco is also used to accompany fiestas and celebrations, such as the famous gipsy weddings, and thus we may find highly characteristic words such as adoro ’I adore’, quiero ’I want/love’, and amor ’love’.
Finally, we find a fourth group involving abstract concepts such as perdón ’forgiveness’, alma ’soul’ or libertad ’freedom’ in palos like soleá, and religious references to God, Christ and to some saints such as Santiago ’Saint James’, Santa Ana ’Saint Anne’, or the Virgin Mary.
@@
’Cadiz is not called Cadiz
@@
in which the words relicario ’reliquary’ or patrona ’patron saint’ acquire relevance without being semantically connected to the topics of the rest of the lyrics.
@@
Among the essential words of alegrías genre, we find terms related to geography such as Muralla Real ’Royal Wall’, Navarra ’Navarre’, Cádiz ’Cadiz’, San Fernando ’Saint Ferdinand’, mar ’sea’, barrio ’neighborhood’, and Zaragoza, the capital of Aragon. This is consistent with the purported origin of alegrías stemming from the dynamization of Aragonese jota music during the French occupation, with a clear Aragonese influence ([14] p. 325-236).
@@
Bulerías are one of the most gipsy and festive palos. With origins in festivity and mockery [42], bulerías often feature improvisations at the end of musical performances attended by the gipsy community or in gipsy neighborhoods of Spanish cities. In bulerías we find multiple stanzas regarding the virtues of being a gipsy, featuring anecdotes about beauty, kindness, and recurring themes of love, including essential words such as gitano/a ’gypsy (sing.)’, primo/a ’cousin (sing.)’, quiero ’I want/love’, querer ’to love’, contigo ’with you’, bonita ’beautiful (fem. sing.)’. Another interesting set of essential words is that of female names, such as María, Dolores, or Amparo. These are usually Spanish polysemes – María may refer to the Virgin Mary, amparo ’shelter’, dolores ’pains’. The pains, the bad, and the sorrows are typical of bulerías, in conjunction with Seville, family, religion, and love.
@@
In malagueñas and fandangos genres we find references to deep sorrows and religion (llorar ’to cry’, llorando ’crying’, pena ’pain, sorrow’, traición ’betrayal’, muere ’he/she dies’, morir ’to die’; muerte ’death’ is the most characteristic word of malagueñas), but also intensely to the love to women (in fandangos, mujer ’woman’ is the most essential word, followed by quería ’I wanted/I loved’ and querer ’to want/love’), as fandangos are usually devoted to sending love messages or love declarations. An interesting lexical feature of fandangos is the essential phraseme cinco sentidos ’five senses’, which is directly related to love dependence. As the lover constitutes the five senses of the singer, when the lover is lost, the singer is found suffering and losing the five senses. Also among the essential words of malagueñas we find lexical references to desperation and profound emotions, such as amaba ’I loved’ and delirio ’delirium’. Both malagueñas and fandangos are referred to as the cantes ’songs’ of Málaga, so it is not surprising that they cover similar lexical fields.
@@
Seguiriyas are believed to be, along with soleá, the gipsiest and purest Flamenco styles. Indeed, both styles include characteristic references to Caló lexicon, such as duquelas ’fatigues, pains’, camelar ’to desire, consent, enamor’, Undebel ’God’, and terelar ’to have, possess’. Seguiriyas and soleá also primarily refer to situations of distress and suffering. For this reason, we can find references to sad, deep concepts such as Dios ’God’, culpa ’guilt’, noche ’night’, ciencia ’science’, infierno ’hell’, muerte ’death’, alma ’soul’, tierra ’earth’, madre ’mother’, or pena ’pain, sorrow’. Interestingly, seguiriyas feature specific references to clavo ’clove’, clavito ’little clove’, canela ’cinnamon’, huele ’smells’, and hueles ’you smell’, all of them related to a formulaic expression related to a bath ritual believed in the Romani culture to have beneficial effects in life.

### v2

@@ I Introduction
Flamenco, a rich oral musical tradition originating in Andalusia, a region in southern Spain, reflects a diverse tapestry of cultural influences. Throughout the centuries, this region, along with its musical expression, has been significantly shaped by various cultural settlements. Notably, the influences of Jews, Muslims, Christians, and folkloric music from different Spanish regions are discernible contributors, yet undeniably, the cultural imprint of Andalusian gipsies is deeply embedded in the essence of Flamenco ([18], p. 469-473). Recognizing its distinct features and profound significance for the cultural identity of Andalusia, Flamenco was honored with inclusion in the UNESCO List of Intangible Cultural Heritage of Humanity in 2010 [51].
@@
Flamenco, as an oral music tradition, involves the transmission of songs and musical nuances across generations. Consequently, written scores are a rarity, given the intricate and complex melodic ornamentation inherent in vocal melodies. Even the records of renowned Flamenco festivals lack information about some features of the performances, including the lyrics [19]. The process of manual annotation, essential for capturing these subtleties, is not only exceedingly time-consuming but also inevitably involves a subjective layer of interpretation.
@@
To analyze large volumes of text, we use Natural Language Processing (NLP) tools. The newborn field of Lyrics Information Processing (LIP) [52] has proven to be successful in retrieving all kinds of information through the use of NLP for a computational analysis of lyrics outside Flamenco. Related works on automatic analysis of lyrics cover a wide range of applications such as the author classification of German songs [34], automatic detection of sexism in English songs [7], the age classification of Indonesian lyrics [47] or the identification of different writing styles in Taylor Swift’s songs [25], among others [30, 15, 2, 39]. However, to our knowledge, the computational analysis of Flamenco singers’ production is limited to studies involving analysis of melody and related features [20, 24] rather than the content of their lyrics. We thus face an absence of LIP applications for Flamenco lyrics analysis.
@@
    Can Flamenco songs be properly assigned to their corresponding palo solely by means of their lexicon? If so, what is the characteristic lexicon of each palo?
@@ II.1 Dataset
We analyze a corpus of Flamenco lyrics curated by Norman P. Kliman on his website [27]. This corpus contains 3380 lyrics spanning 78 different styles, or palos. The lyrics were transcribed by both dedicated aficionados and Kliman himself. Metadata such as author, album, year, and palo was not assigned by the transcribers, but rather reflects the information provided by the original record label or other publishing sources. As mentioned in Sec. I, palo cathegorization is subject of melodical and musical interpretations, which can vary depending on the listener and the tradition. Importantly, we use the palo metadata provided by the record labels of the albums as the ground truth for the supervised learning algorithm (see Sec. II.2 and Appendix B.1). The corpus features 276822 tokens (words) and 12981 types (unique words). The corpus is varied and diverse, exhibiting a structure characteristic of natural language corpora (see Appendix A.1).
@@
Also, for the sake of clarity while reading this manuscript, we specify in Table 1 the acronyms with which we will refer to the aforementioned styles throughout the text and in the Figures. It is important to note that categorizing Flamenco songs into styles by ear can be subjective, but for validation purposes, we rely on the classification provided by the album in which each song appears. We consider this classification as the ground truth.
@@
After preprocessing the corpus, a necessary step previous to their computational analysis which involves deleting punctuation signs and accents and distinguishing between common and proper nouns (see Appendix A.2), we can extract information from the lexical structure of the lyrics. We now proceed to describe several findings in the process of corpus characterization.
@@
One way to quantify the lexical richness of a text is to measure its vocabulary size $|V|$, i.e., the amount of types. In Fig. 2(a) we show the vocabulary size and the amount of tokens of the eight different styles or palos, computed as the vocabulary size of an aggregation of all the lyrics of each palo. Bulerías account for the highest number of types used while malagueñas restrict the range of use to a vocabulary almost 6 times lower in size. The overall ranking of token and type frequencies remains consistent across genres, with the exception of seguiriyas.
In the context of song-wise analysis, we may also observe in Fig. 2(b) that bulerías have the widest distribution of types used among songs, while malagueñas, fandangos and seguiriyas are restricted to a similar number of types among songs. Cases like malagueñas or alegrías resemble bimodal distributions, indicating that mainly two different vocabulary sizes characterize their lyrics. With these remarks, we can already infer that bulerías serve as the most polyvalent style, with the broadest vocabulary size distribution and the highest amount of types.
Figure 2: (a) Amount of tokens $L$ (bars on the right) and types $|V|$ (bars on the left) for each style, and (b) distribution of the vocabulary size, $|V|$ (number of types) for each style. The continuous distribution is derived from a discrete histogram using the kernel density estimation method [11].
@@
In our case, Eq. 1 is affected by the fact that some genres use more types than others, as we see in Fig. 2(a), while also having different amount of tokens and lyrics. For that reason, we use the standardized $TTR$, sTTRsTTR, which is defined as the $TTR$ averaged over several portions of the same document, with all the portions having the same number of tokens. For each palo we create a document compiling all the genre’s lyrics. Malagueñas is the genre with the least amount of tokens. Then, we take 50 portions of each genre with the length of the entire malagueñas subcorpus and compute the mean $sTTR$ and its error except for malagueñas, for which we take the entire genre a single time without error. We have checked that increasing the portion number above 50 does not significantly alter our results. We show the results in Fig. 3, where the solid line accounts for the $sTTR$ of the entire corpus, our null model for comparison.
@@
The different values of $sTTR$ provide a partial answer to RQ1 by revealing that various Flamenco genres are distinguished by their unique vocabulary distributions and levels of lexical richness. However, it is not clear that these measurements would allow for an unambiguous identification of a palo given some lyrics. For example, the vocabulary distribution is heavily influenced by the specific corpus used, and the reliability of $sTTR$ as a measure of lexical diversity independent of text length has been challenged [31]. Moreover, $sTTR$ values for some palos are close enough to be indistinguishable, as in the case of tangos and tientos, or bulerías and alegrías. To provide a comprehensive answer to RQ1, we need a more sophisticated approach. Thus, we will now examine the specific lexicon utilized by each palo using a machine learning algorithm which has proven to be useful in text analysis.
@@ II.2 Palo classification
We employ a Multinomial Naive Bayes (MNB) model [26], a machine learning classification algorithm, to classify lyrics into distinct Flamenco genres or palos solely based on their unigram lexical content. Initially, the corpus is partitioned into a training set comprising 85% of the data and a validation set containing the remaining 15%. The algorithm is then trained using the training set, considering the palo labels provided in the dataset as the ground truth. Subsequently, the model performance is assessed using the validation set, wherein the genres assigned by the model are compared against the true genre labels associated with the lyrics. Further details about how this classification algorithm works are included in Appendix B.1.
@@
Quite generally, different arrangements of lyrics for the training and validation sets yield different results. For this reason, we compute the accuracy of the model (the ratio of correct assignments) for 100 different trainings, obtaining the results shown in Fig. 5(a).
Figure 5: (a) Violin plot of the distribution of accuracies obtained for each genre across 100 trainings of the Multinomial Naive Bayes model using various corpus splittings for training and validation, and (b) average confusion matrix of the MNB for 100 different trainings, showing the percentage of confusions of each Flamenco style with other genres. The normalization is based only on the confusions.
Despite the fact that genres like soleá (85% mean accuracy), alegrías (88%) and seguiriyas (91%) are classified with outstanding accuracies, a few lyrics are incorrectly classified and therefore confused by the model. To examine these discrepancies, in Fig. 5(b) we plot the average confusion matrix of 100 different trainings and account only for the incorrect assignments.
@@
It is evident how the aforementioned confusions affect the accuracy of the model, as in the cases of tangos (mean accuracy of 50%) and tientos (61%). However, Fig. 5(a) shows that soleá, alegrías and seguiriyas are classified with outstanding accuracies. Thus, we can state that a classification of Flamenco lyrics into their corresponding palos based solely on their lexicon is quantitatively possible (RQ1), and that the inaccuracies are due to interesting palo relationships. Now, let us extract the lexical features that allow the model to classify the lyrics, i.e., the characteristic lexicon of each style.
@@ II.3 Characteristic lexicon extraction
The MNB computes a metric called log-probability, $\log P(w|C)$, of each word $w$ for each palo $C$ (see Appendix B.1). The higher $P(w|C)$ is, the more characteristic $w$ is for $C$. We compute $P(w|C)$ for several trainings of the algorithm and obtain results as shown in Fig. 6, where we can see the 15 most characteristic words, i.e., the 15 words with the largest value of $\log P(w|C)$, for two example styles, $C=\mathrm{A}$ (alegrías) [Fig. 6(a)] and $C=\mathrm{Se}$ (seguiriyas) [Fig. 6(c)].
Figure 6: Summary of the most characteristic words for alegrías and seguiriyas. We plot the log-probability of the 15 most characteristoc words for (a) alegrías and (c) seguiriyas. Additionally, we show two wordcloud diagrams in which the words’ sizes reflect how characteristic they are for (b) alegrías and (d) seguiriyas.
While alegrías feature as most characteristic words references to places such as Cadiz, Muralla Real ‘Royal Wall’ or mar ’sea’ (all related to Cadiz and its harbors), and other less specific words such as llevar ‘to take/bring’, voy ‘I go’, conmigo ‘with me’, seguiriyas are biased towards abstract and profound concepts such as Dios ‘God’, alma ‘soul’, libertad ‘freedom’ and desgracia ‘misfortune’. We will later interpret the characteristic words in Sec. III, and see how these words clearly differentiate between genres.
@@
The number of essential words within each palo serves for distinguishing among various Flamenco genres. Table 2 presents the count of essential words per palo, denoted as $N_{e}$. Notably, while malagueñas have 16 essential words, soleá’s essential lexicon is significantly larger, averaging 186 words, i.e., nearly 12 times more. Despite this, as illustrated in Fig. 2(a), malagueñas exhibit a far smaller vocabulary compared to soleá. To account for the impact of vocabulary size on the essential lexicon, we normalize $N_{e}$ by the vocabulary size, $|V|$, as depicted in Fig. 8. When adjusted for vocabulary size, seguiriyas boasts the largest essential lexicon, while soleá’s normalized essential lexicon (5.5%) is only four times greater than that of malagueñas (1.5%). The complete lists of these words for all Flamenco genres in descending order of log-probability are included in Appendix C. We emphasize that the relevance of these words is automatically detected and is therefore free of subjective biases. For example, it also identifies certain amount of hapax legomena at palo level as significant, as discussed in Appendix B.3. This is a clear advantage of our application of computational classification methods.
@@ II.4 Relationship between palos
The distances between Flamenco styles are therefore helpful in addressing RQ2. To provide a more concrete analysis, we adopt a network representation. In Fig. 10(a), we present a fully-connected network illustrating the distances between styles, enabling a straightforward visual assessment of their relationships. Each node represents a style, and the edges vary in thickness and color depending on the cosine distance values. The network structure provides insights into the relationships among different palos in several ways. As observed in the dendrogram, alegrías emerges as the most lexically distinct palo. The remaining palos form a more tightly interconnected network, with bulerías, soleá, and tangos prominently contributing to its connectivity. Additionally, the size of the nodes in the network graph corresponds to their closeness centrality, a measure in network analysis that quantifies how close a node is to all other nodes in the network ([14], p. 146), suggesting their significance in facilitating connections between different musical forms. This network representation proves advantageous as it facilitates the extraction of spanning trees ([14], p. 27), which are sub-networks that connect all genres such that each genre is linked directly or indirectly to every other genre, without creating any cycles, thereby maintaining the network connectivity. By selecting the minimum spanning tree (MST), which minimizes the sum of distances between palos, we obtain the representation depicted in Fig. 10(b), i.e., the MST of Flamenco palos. This approach enables a structured exploration of the relationships between palos, shedding light on their hierarchical organization within the Flamenco genre.
Figure 10: (a) Network relationships between different styles or palos, with each node representing a palo or style. The thickness and color of the edges indicate the distances between the palos, and the size of each node is proportional to its closeness centrality. (b) Minimum spanning tree of Flamenco palos, showing four branches connected by bulerías. The labels on the nodes follow the acronyms of Table 1.
@@ III Discussion
At first glance, regarding the $sTTR$ analysis, in Fig. 3 we can see that $sTTR$ is genre-distinctive. Indeed, returning to the results of the vocabulary size distribution in Fig. 2(b), if we focus on the $sTTR$ values, we can see how bulerías make lower use of their types than malagueñas do, even though the vocabulary of the latter is significantly lower than the vocabulary of the former. This indicates that while bulerías are rather varied and move within a wide range of topics, this genre makes relatively less use of their topics than malagueñas do. Values above the null model’s $sTTR$ indicate a more developed use of a specific set of types, while values below indicate either a wider use of types or a less developed use of the available types.
@@
First, we find words related to geographical places and day-to-day life. As previously mentioned, Flamenco lyrics are transmitted orally and encompass a tradition of livings and history. Thus it is common to find references to these topics. Thus, while alegrías feature lexicon related to Cádiz, Navarre and titirimundis ‘puppets show’, malagueñas refer to the Corte ‘Court, Capital’ (Madrid), tangos talk about wars and bulerías feature references to Seville and proper names like Amparo, Dolores or María.
Secondly, we also find references to gipsies and their families and culture, due to the intense influence of this ethnic group in Flamenco art throughout its history. We obtain specific words like gitano/a ‘gipsy’, and common words like madre ‘mother’, padre ‘father’, or primo/a ‘cousin’, usually in a gipsy context. Interestingly, seguiriyas and soleá genres even feature Caló vocabulary [42] as characteristic words. The use of Caló, a dialect of Romani language spoken by the gipsies of Spain ([33], p. 125), further connects these genres to the gipsy people.
We find evidence for the belief that Flamenco accompanied gipsies and Andalusian people in times of sorrow and high suffering [22] in lexicon related to deep pain, such as dolor ‘pain’, desgracia ‘misfortune’, llorar ‘to cry’, pena ‘sorrow’, traición ‘betrayal’, muerta ‘dead (singular feminine)’. This semantic field constitutes a third group, along with lexicon related to joy and love. Flamenco is also used to accompany fiestas and celebrations, such as the famous gipsy weddings, and thus we may find highly characteristic words such as adoro ‘I adore’, quiero ‘I want/love’, and amor ‘love’.
Finally, we find a fourth group involving abstract concepts such as perdón ‘forgiveness’, alma ‘soul’ or libertad ‘freedom’ in palos like soleá, and religious references to God, Christ and to some saints such as Santiago ‘Saint James’, Santa Ana ‘Saint Anne’, or the Virgin Mary.
@@
‘Cadiz is not called Cadiz
@@
in which the words relicario ‘reliquary’ or patrona ‘patron saint’ acquire relevance without being semantically connected to the topics of the rest of the lyrics.
@@
Among the essential words of alegrías genre, we find terms related to geography such as Muralla Real ‘Royal Wall’, Navarra ‘Navarre’, Cádiz ‘Cadiz’, San Fernando ‘Saint Ferdinand’, mar ‘sea’, barrio ‘neighborhood’, and Zaragoza, the capital of Aragon. This is consistent with the purported origin of alegrías stemming from the dynamization of Aragonese jota music during the French occupation, with a clear Aragonese influence ([18] p. 325-236).
@@
Bulerías are one of the most gipsy and festive palos. With origins in festivity and mockery [9], bulerías often feature improvisations at the end of musical performances attended by the gipsy community or in gipsy neighborhoods of Spanish cities. In bulerías we find multiple stanzas regarding the virtues of being a gipsy, featuring anecdotes about beauty, kindness, and recurring themes of love, including essential words such as gitano/a ‘gypsy (sing.)’, primo/a ‘cousin (sing.)’, quiero ‘I want/love’, querer ‘to love’, contigo ‘with you’, bonita ‘beautiful (fem. sing.)’. Another interesting set of essential words is that of female names, such as María, Dolores, or Amparo. These are usually Spanish polysemes – María may refer to the Virgin Mary, amparo ‘shelter’, dolores ‘pains’. The pains, the bad, and the sorrows are typical of bulerías, in conjunction with Seville, family, religion, and love.
@@
In malagueñas and fandangos genres we find references to deep sorrows and religion (llorar ‘to cry’, llorando ‘crying’, pena ‘pain, sorrow’, traición ‘betrayal’, muere ‘he/she dies’, morir ‘to die’; muerte ‘death’ is the most characteristic word of malagueñas), but also intensely to the love to women (in fandangos, mujer ‘woman’ is the most essential word, followed by quería ‘I wanted/I loved’ and querer ‘to want/love’), as fandangos are usually devoted to sending love messages or love declarations. An interesting lexical feature of fandangos is the essential phraseme cinco sentidos ‘five senses’, which is directly related to love dependence. As the lover constitutes the five senses of the singer, when the lover is lost, the singer is found suffering and losing the five senses. Also among the essential words of malagueñas we find lexical references to desperation and profound emotions, such as amaba ‘I loved’ and delirio ‘delirium’. Both malagueñas and fandangos are referred to as the cantes ‘songs’ of Málaga, so it is not surprising that they cover similar lexical fields.
@@
Seguiriyas are believed to be, along with soleá, the gipsiest and purest Flamenco styles. Indeed, both styles include characteristic references to Caló lexicon, such as duquelas ‘fatigues, pains’, camelar ‘to desire, consent, enamor’, Undebel ‘God’, and terelar ‘to have, possess’. Seguiriyas and soleá also primarily refer to situations of distress and suffering. For this reason, we can find references to sad, deep concepts such as Dios ‘God’, culpa ‘guilt’, noche ‘night’, ciencia ‘science’, infierno ‘hell’, muerte ‘death’, alma ‘soul’, tierra ‘earth’, madre ‘mother’, or pena ‘pain, sorrow’. Interestingly, seguiriyas feature specific references to clavo ‘clove’, clavito ‘little clove’, canela ‘cinnamon’, huele ‘smells’, and hueles ‘you smell’, all of them related to a formulaic expression related to a bath ritual believed in the Romani culture to have beneficial effects in life.